Evaluate Integrals. Show complete solution.

3. Integral of 1/ 2u^2 minus 3u^3/2 minus 2u^5/ 3 du
4. Integral of 2t^6 + 5t^2 +7 / t^2 dt
1
Expert's answer
2019-11-11T10:29:42-0500

"3) \int" (1/(2u2) -3u3/2-2u5/3)du= (-1/2u)-3u4/8-2u6/6+C =(-1/2u)-3u4/8-u6/3+C.

4) "\int" (2t6+5t2+7/t2)dt= 2t7/7+5t3/3-7/t +C.